OpenStudy (anonymous):

solve for the ff. exponentials: (3^2x)(2^1-x)=5^x 5^x + 24= 25(5^-x) log2(logx4)=1

OpenStudy (anonymous):

what is ff

OpenStudy (anonymous):

(3^2x)(2^(1-x))=5^x or (3^2x)(2^(1)-x)=5^x which of these do you mean

jhonyy9 (jhonyy9):

the last one will be log2(logx4)=1 there are log2 what signe logarithm base 2 ? and logx4 signe logarithm base x of 4 ?

OpenStudy (wwe123):

for the ff ??

jhonyy9 (jhonyy9):

can you please clarifie this ,,ff" what mean you with this ? and there are 2x and 1-x and for 5 the x like exponents all ?

OpenStudy (anonymous):

ff. means following. im sorry :))

jhonyy9 (jhonyy9):

ok so can you answer my questions from exponents and please us parantheses for clarifie this thank you